OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3) Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3) @Mathematics

OpenStudy (sasogeek):

4(x-6) = -2(x-3) 4x-24 =-2x+6 4x+2x=6+24 6x=30 x=5

OpenStudy (anonymous):

4(x-6) = -2(x-3) 4x-24 = -2x +6 +2x +24 +2x +24 6x = 30 x = 5